# Polymatroid Schubert varieties

### v1

@@ 4. Polymatroid Schubert varieties
Let $T=\{i:\overline{s}_{i}=n_{i}\}$. This set is nonempty: if we suppose falsely that $T=\emptyset$, then by 3.2, $\mathbf{s}=\overline{\mathbf{s}}$, hence $\operatorname{rk}_{P}(\mathbf{s})=|\mathbf{s}|=d$, contrary to the fact that $\mathbf{s}$ is not a basis. By 3.2,
$\dim\pi_{T}(Y_{V})=\operatorname{rk}_{P}(T)=\operatorname{rk}_{P,\mathbf{n}}(\overline{\mathbf{s}}^{\mathrm{geo}})=\operatorname{rk}_{P,\mathbf{n}}(\overline{\mathbf{s}})-\sum_{i\in E\setminus T}s_{i}<d-\sum_{i\in E\setminus T}s_{i}=\sum_{i\in T}s_{i}.$
Consequently, if $\mathbf{s}^{\prime}=\mathbf{s}-\sum_{i\in E\setminus T}s_{i}\mathbf{e}_{i}$, then
$c_{\mathbf{s}}[\mathrm{pt}]=[\mathbb{P}^{\mathbf{n}-\mathbf{s}}]\cdot[Y_{V}]=[\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot([\mathbb{P}^{\mathbf{n}-\mathbf{s}^{\prime}}]\cdot[Y_{V}])=[\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot 0=0.$
@@
The Chow class of $Y_{V}$ in $A_{d}(\mathbb{P}^{\mathbf{n}})$ is of the form
$[Y_{V}]=\sum_{\mathbf{b}\text{ basis of $P$}}c_{\mathbf{b}}\,[\mathbb{P}^{\mathbf{b}}],$
with each coefficient $c_{\mathbf{b}}>0$.
@@
Let $I:\mathbb{K}^{\mathbf{n}}\to O_{\mathbf{n}}\cong\mathbb{P}^{\mathbf{n}}$ be the usual map from $\prod_{i=1}^{N}\mathbb{K}^{n_{i}}$ to $\prod_{i=1}^{N}\mathbb{P}^{n_{i}}$, given by $\frac{b_{ij}}{b_{i0}}(I(\mathbf{a}))=a_{ij}$. Write $\overline{V}$ for the closure of $I(V)$ in $\mathbb{P}^{\mathbf{n}}$. The Chow class of $\overline{V}$ was computed in [Li18], and is given by
@@
$\frac{b_{i\ell}}{b_{i0}}(\iota^{\prime}_{\circ}(\mathbf{a},t))=\sum_{j_{1}+2j_{2}+\cdots+kj_{k}=\ell}t^{k-1}\frac{a_{i,1}^{j_{1}}a_{i,2}^{j_{2}}\cdots a_{i,k}^{j_{k}}}{j_{1}!j_{2}!\cdots j_{k}!},\quad 1\leq\ell\leq n_{i},$
@@
Let $X$ be the closure of $\iota^{\prime}(V\times\mathbb{A}^{1})$ in $\mathbb{P}^{\mathbf{n}}\times\mathbb{A}^{1}$. Observe that $X_{0}\supset\overline{V}$. On the other hand, there is a diagram
@@
The polymatroid Schubert variety $Y_{V}$ is $V$-equivariant, hence is a union of $V$-orbits. To guarantee that the number of orbits is finite, we require $V$ to be polymatroid general (henceforth, p.g.), meaning that for each multiset $\mathbf{s}\leq\mathbf{n}$,
$\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})=\operatorname{codim}_{V}V\cap S_{\mathbf{s}}.$
@@
so $\operatorname{codim}_{V^{\prime}}V^{\prime}\cap S_{\mathbf{s}^{\prime}}=\operatorname{codim}_{V}V\cap S_{\mathbf{s}}=\operatorname{rk}_{P}(\mathbf{s})$. The multisets $\mathbf{s}$ and $\mathbf{s}^{\prime}$ are represented by the same set $S\subset\widetilde{E}\setminus E_{i}$, so
$\operatorname{codim}_{V^{\prime}}V^{\prime}\cap S_{\mathbf{s}^{\prime}}=\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})=\operatorname{rk}_{\widetilde{P}}(S)=\operatorname{rk}_{\widetilde{P}\setminus E_{i}}(S)=\operatorname{rk}_{\widetilde{P\setminus i}}(S)=\operatorname{rk}_{P\setminus i,\pi_{E\setminus i}(\mathbf{n})}(\mathbf{s}^{\prime}),$
@@
Suppose $V\subset\mathbb{K}^{\mathbf{n}}$ is a p.g. subspace and $A\subset E$. If $H$ is the preimage in $\mathbb{K}^{\mathbf{n}}$ of a general [^6] By which we mean a hyperplane such that $H\cap V$ realizes $T_{A}P$. See also 2.15. hyperplane in $\mathbb{K}^{\pi_{A}(\mathbf{n})}$, then $V\cap H$ is p.g.
@@
$\operatorname{codim}_{V\cap H}V\cap H\cap S_{\mathbf{s}}=\begin{cases}\operatorname{codim}_{V}V\cap S_{\mathbf{s}}-1,&\text{if $V\cap S_{\mathbf{s}}\subset V\cap H$}\\ \operatorname{codim}_{V}V\cap S_{\mathbf{s}},&\text{otherwise}\end{cases}=\begin{cases}\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})-1,&\text{if $V\cap S_{\mathbf{s}}\subset V\cap H$}\\ \operatorname{rk}_{P,\mathbf{n}}(\mathbf{s}),&\text{otherwise.}\end{cases}$
To finish, observe that $A\subset\{i:\overline{s}_{i}=n_{i}\}$ if and only if $V\cap S_{\mathbf{s}}\subset V\cap H$ for a Zariski-general $H$. ∎
@@
$\dim(V\cdot x)=\operatorname{codim}_{V}(V\cap S_{\mathbf{s}})=\operatorname{rk}_{P}(\mathbf{s}).$ (*)
@@
Otherwise, suppose $N>1$. If $\operatorname{rk}(\mathbf{s})=d$, then the dimension formula implies $Y_{V}\cap O_{\mathbf{s}}$ is nonempty if and only if $\mathbf{s}=\mathbf{n}$. Otherwise, if $\operatorname{rk}\mathbf{s}<d$, then there is $i\in E$ such that $\overline{s}_{i}<n_{i}$. The projection $Y_{V}\to Y_{\pi_{E\setminus i}(V)}$ sends $Y_{V}\cap O_{\mathbf{s}}$ into $Y_{\pi_{E\setminus i}(V)}\cap O_{\pi_{E\setminus i}(\mathbf{s})}$. By 3.10, $\pi_{E\setminus i}(\mathbf{s})$ is not a combinatorial flat of $\pi_{E\setminus i}(V)$, so $Y_{\pi_{E\setminus i}(V)}\cap O_{\pi_{E\setminus i}(\mathbf{s})}$ is empty by induction on $N$. This means $Y_{V}\cap O_{\mathbf{s}}$ is empty too. ∎
@@
We complete the proof for proper combinatorial flats by inducting on $N$. First consider $N=1$, so that $Y_{V}\subset\mathbb{P}^{n_{1}}$. By 4.9, $Y_{V}\cap\mathbb{P}^{n_{1}-1}\subset\mathbb{P}^{\dim V-1}$. Since $Y_{V}\cap\mathbb{P}^{n_{1}-1}$ is a divisor on $Y_{V}$, we must have $Y_{V}\cap\mathbb{P}^{n_{1}-1}=\mathbb{P}^{\dim V-1}$. Since $V$ is polymatroid general, $O_{i}\subset\mathbb{P}^{\dim V-1}$ is a single $V$-orbit, and $\overline{Y_{V}\cap O_{i}}=\mathbb{P}^{i}$ as desired.
@@
Otherwise, suppose $\dim V>1$. For a general hyperplane $H$ in $\mathbb{K}^{\mathbf{n}}$, $Y_{V\cap H}\cap O_{\mathbf{s}}$ is nonempty for every combinatorial flat $\mathbf{s}$ of corank greater than 1 by the induction hypothesis and 3.11. By the same argument, if $F=\{i:s_{i}=n_{i}\}$ is nonempty, then $Y_{V\cap H}\cap O_{\mathbf{s}}$ is nonempty for $H\subset\mathbb{K}^{\mathbf{n}}$ the preimage of a general hyperplane in $\mathbb{K}^{\pi_{F}(\mathbf{n})}$.
@@
where $C$ runs over all irreducible components of $Y_{V}\cap H$. By 4.9, $Y_{V}\cap H\subset\cup_{\mathbf{s}^{\prime}}Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}$, where the union runs over all corank 1 combinatorial flats $\mathbf{s}^{\prime}\leq\mathbf{n}-\mathbf{e}_{i}$. By 4.10, each irreducible component $C$ of $Y_{V}\cap H$ is among those of $\cup_{\mathbf{s}^{\prime}}Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}$, so
@@
with $m_{\mathbf{s}^{\prime}}=0$ if $\dim Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}<d-1$. By 4.10 and 4.6, we may further expand as
@@
Characterization of when $Y_{V}$ intersects $O_{\mathbf{s}}$ is 4.9 and 4.11. 1.3i follows from 4.10, and 1.3iii is obtained by combining 4.10 and 4.6.
@@
Fix $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$. By 3.10, $\mathbf{s}^{\prime}<\mathbf{s}$ is a combinatorial flat of $V$ if and only if $\mathbf{s}_{i}^{\prime}\leq\mathbf{s}_{i}$ for all $i$ and $(\mathbf{s}_{j}^{\prime})_{j:s_{j}=n_{j}}$ is a combinatorial flat of $V^{\prime}:=\pi_{\{j:s_{j}=n_{j}\}}(V)$. Such multisets $\mathbf{s}^{\prime}$ are precisely those that index the nonempty strata of $\overline{Y_{V}\cap O_{\mathbf{s}}}$, since $\overline{Y_{V}\cap O_{\mathbf{s}}}=Y_{V^{\prime}}\times\prod_{s_{i}<n_{i}}\mathbb{P}^{s_{i}}$ by 4.10. 4.10 also implies $O_{\mathbf{s}^{\prime}}\cap(Y_{V^{\prime}}\times\prod_{s_{i}<n_{i}}\mathbb{P}^{s_{i}})=O_{\mathbf{s}^{\prime}}\cap Y_{V}$ whenever both sides are nonempty, which completes the proof. ∎
@@
Theorems 1.3 describes properties of polymatroid Schubert varieties that are well-known for the matroid Schubert varieties of [AB16], which correspond to the case $\mathbf{n}=(1,\ldots,1)$. One further such property follows: when $\mathbf{n}=(1,\ldots,1)$, it is well-known that changing $V$ by rescaling each coordinate of $\mathbb{K}^{\mathbf{n}}=\mathbb{K}^{N}$ does not change the isomorphism class of $Y_{V}$. For general $\mathbf{n}$, the isomorphism class of $Y_{V}$ is unchanged if we perform a weighted rescaling of each factor $\mathbb{K}^{n_{i}}$, as described in Section 4.2.
@@
Assume that $V$ realizes a simple polymatroid. The subspaces $V_{ij}=V\cap S_{j\mathbf{e}_{i}}$ for $1\leq i\leq N$ and $1\leq j\leq n_{i}$ comprise a collection of $N$ flags in $V$. The combinatorial flats record intersection pattern of these flags, as $\operatorname{rk}(\mathbf{s})=\operatorname{codim}_{V}V_{1,s_{1}}\cap V_{2,s_{2}}\cap\cdots\cap V_{N,s_{N}}$. The “polymatroid general” condition on $V$ guarantees that these flags are in general position with respect to one another (see also 3.7).
@@
In this section, we establish 1.4, which gives a formula for the cohomology ring of $Y_{V}$. Throughout, we assume that $V$ is polymatroid general. For an independent multiset $\mathbf{b}$, let $c_{\mathbf{b}}>0$ be the coefficient of $[\mathbb{P}^{\mathbf{b}}]$ in $[Y_{V}\cap\mathbb{P}^{\overline{\mathbf{b}}}]$, as in the statement of 1.3.
@@
Let $\kappa:Y_{V}\to\mathbb{P}^{\mathbf{n}}$ denote the inclusion. By 1.3iii, the classes $[Y_{V}\cap\mathbb{P}^{\mathbf{s}}]$ all have pairwise disjoint supports, so the pushforward $A_{*}Y_{V}\to A_{*}\mathbb{P}^{\mathbf{n}}$ is injective.
The sets $O_{\mathbf{s}}$ and $Y_{V}\cap O_{\mathbf{s}}$ comprise cell decompositions of $\mathbb{P}^{\mathbf{n}}$ and $Y_{V}$, respectively. Hence, the Chow groups of these varieties are naturally isomorphic to their Borel-Moore homology groups via the cycle class map [Ful98, Example 19.1.11]. Since $\mathbb{P}^{\mathbf{n}}$ and $Y_{V}$ are compact, their Borel-Moore homology is equal to their singular homology. Summing up, we’ve learned that the singular homology groups of $Y_{V}$ are all free and that the homology pushforward $\kappa_{*}:H_{*}(Y_{V};{\mathbb{Z}})\to H_{*}(\mathbb{P}^{\mathbf{n}};{\mathbb{Z}})$ is injective. Applying the Universal Coefficient Theorem, we learn that $\kappa^{*}:H^{*}(\mathbb{P}^{\mathbf{n}};\mathbb{Q})\to H^{*}(Y_{V};\mathbb{Q})$ is surjective.

### v2

@@ 4. Schubert varieties of subspace arrangements
Let $T=\{i:\overline{s}_{i}=n_{i}\}$. This set is nonempty: if we suppose falsely that $T=\emptyset$, then by 3.2, $\mathbf{s}=\overline{\mathbf{s}}$. This means that $\mathbf{s}$ is a combinatorial flat satisfying $\mathbf{s}^{\mathrm{geo}}=\mathbf{0}$, so by 3.2 once more,
$\operatorname{rk}_{P}(\mathbf{s})=\operatorname{rk}_{P}(\mathbf{s}^{\mathrm{geo}})+|\mathbf{s}-\mathbf{s}^{\mathrm{geo}}|=0+|\mathbf{s}|=d,$
contrary to the fact that $\mathbf{s}$ is not a basis.
Let $\mathbf{s}^{\prime}=\mathbf{s}-\sum_{i\in E\setminus T}s_{i}\mathbf{e}_{i}$. The pushforward $\pi_{T*}:A_{0}(\mathbb{P}^{\mathbf{n}})\to A_{0}(\mathbb{P}^{\pi_{T}(\mathbf{n})})$ is an isomorphism, so $c_{\mathbf{s}}=0$ if and only if $\pi_{T*}(c_{\mathbf{s}}[\mathrm{pt}])$ is zero in $A_{*}(\mathbb{P}^{\pi_{T}(\mathbf{n})})$. By the projection formula,
$\pi_{T*}(c_{\mathbf{s}}[\mathrm{pt}])=\pi_{T*}([\mathbb{P}^{\mathbf{n}-\mathbf{s}}]\cdot[Y_{V}])=\pi_{T*}([\mathbb{P}^{\mathbf{n}-\mathbf{s}^{\prime}}]\cdot[\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot[Y_{V}])=[\mathbb{P}^{\pi_{T}(\mathbf{n}-\mathbf{s}^{\prime})}]\cdot\pi_{T*}([\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot[Y_{V}]).$
The class $[\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot[Y_{V}]$ is represented by a subvariety $Y^{\prime}$ of $Y_{V}$, obtained by intersecting $Y_{V}$ with appropriately general hyperplanes. To show that $[\mathbb{P}^{\pi_{T}(\mathbf{n}-\mathbf{s}^{\prime})}]\cdot\pi_{T*}([\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot[Y_{V}])=[\mathbb{P}^{\pi_{T}(\mathbf{n}-\mathbf{s}^{\prime})}]\cdot\pi_{T*}([Y^{\prime}])$ is zero, it suffices to show that $\dim\pi_{T}(Y^{\prime})<|\pi_{T}(\mathbf{s}^{\prime})|$. In fact, this is the case: by 3.2,
$\dim\pi_{T}(Y^{\prime})\leq\dim\pi_{T}(Y_{V})$ $=\operatorname{rk}_{P}(T)=\operatorname{rk}_{P,\mathbf{n}}(\overline{\mathbf{s}}^{\mathrm{geo}})$
$=\operatorname{rk}_{P,\mathbf{n}}(\overline{\mathbf{s}})-\sum_{i\in E\setminus T}s_{i}<d-\sum_{i\in E\setminus T}s_{i}=\sum_{i\in T}s_{i}=|\pi_{T}(\mathbf{s}^{\prime})|.$
We conclude that $c_{\mathbf{s}}=0$. ∎
@@
For each basis $\mathbf{b}$ of $P$, $c_{\mathbf{b}}>0$.
@@
Let $I:\mathbb{K}^{\mathbf{n}}\to O_{\mathbf{n}}\subset\mathbb{P}^{\mathbf{n}}$ be the usual map from $\prod_{i=1}^{N}\mathbb{K}^{n_{i}}$ to $\prod_{i=1}^{N}\mathbb{P}^{n_{i}}$, given by $\frac{b_{ij}}{b_{i0}}(I(\mathbf{a}))=a_{ij}$. Write $\overline{V}$ for the closure of $I(V)$ in $\mathbb{P}^{\mathbf{n}}$. The Chow class of $\overline{V}$ is (see e.g. [CCRL+20]) given by
@@
$\frac{b_{i\ell}}{b_{i0}}(\iota^{\prime}_{\circ}(\mathbf{a},t))=\sum_{j_{1}+2j_{2}+\cdots+\ell j_{\ell}=\ell}t^{j_{1}+j_{2}+\cdots+j_{\ell}-1}\frac{a_{i,1}^{j_{1}}a_{i,2}^{j_{2}}\cdots a_{i,\ell}^{j_{\ell}}}{j_{1}!j_{2}!\cdots j_{\ell}!},\quad 1\leq\ell\leq n_{i},$
@@
Let $X$ be the closure of $\iota^{\prime}(V\times\mathbb{A}^{1})$ in $\mathbb{P}^{\mathbf{n}}\times\mathbb{A}^{1}$. Equivalently, $X$ is the closure of $\iota^{\prime}_{\circ}(V\times\mathbb{G}_{m})$, so $X$ is a flat family over $\mathbb{A}^{1}$ (e.g. by [Liu06, Proposition 4.3.9]).
Observe that $X_{0}\supset\overline{V}$. On the other hand, there is a diagram
@@
Remark 4.7.
Another proof of 4.6 goes as follows: a projection of $O_{\mathbf{n}}\cap Y_{V}$ onto a given $d$-element subset of the coordinates $\{b_{ij}/b_{i0}\}_{i,j}$ is dominant if and only if its induced map on tangent spaces is surjective at a general point of $O_{\mathbf{n}}\cap Y_{V}$. Consequently, to prove that $c_{\mathbf{b}}>0$, it suffices to produce a point $p\in O_{\mathbf{n}}\cap Y_{V}$ and a $d$-element set of coordinates representing $\mathbf{b}$ such that the derivative of the corresponding projection is surjective at $p$.
The derivative of $\iota$ at $0\in V$ identifies $V$ with the tangent space of $Y_{V}$ at the point $p=[1:0:0:\cdots:0]\times\cdots\times[1:0:0:\cdots:0]$. Therefore, if $\mathbf{b}$ is a basis of the polymatroid $P$ associated to $V$, then a $d$-element set of the type desired exists, proving that $c_{\mathbf{b}}>0$.
@@
The Schubert variety $Y_{V}$ is $V$-equivariant, hence is a union of $V$-orbits. To guarantee that the number of orbits is finite, we require $V$ to be polymatroid general (henceforth, p.g.), meaning that for each multiset $\mathbf{s}\leq\mathbf{n}$,
$\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})=\operatorname{codim}_{V}V\cap\operatorname{St}_{\mathbf{s}}.$
@@
so $\operatorname{codim}_{V^{\prime}}V^{\prime}\cap\operatorname{St}_{\mathbf{s}^{\prime}}=\operatorname{codim}_{V}V\cap\operatorname{St}_{\mathbf{s}}=\operatorname{rk}_{P}(\mathbf{s})$. The multisets $\mathbf{s}$ and $\mathbf{s}^{\prime}$ are represented by the same set $S\subset\widetilde{E}\setminus E_{i}$, so
$\operatorname{codim}_{V^{\prime}}V^{\prime}\cap\operatorname{St}_{\mathbf{s}^{\prime}}=\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})=\operatorname{rk}_{\widetilde{P}}(S)=\operatorname{rk}_{\widetilde{P}\setminus E_{i}}(S)=\operatorname{rk}_{\widetilde{P\setminus i}}(S)=\operatorname{rk}_{P\setminus i,\pi_{E\setminus i}(\mathbf{n})}(\mathbf{s}^{\prime}),$
@@
Suppose $V\subset\mathbb{K}^{\mathbf{n}}$ is a p.g. subspace and $A\subset E$. If $H$ is the preimage in $\mathbb{K}^{\mathbf{n}}$ of a general [^6] By which we mean a hyperplane such that $H\cap V$ realizes $T_{A}P$. See also 2.16. hyperplane in $\mathbb{K}^{\pi_{A}(\mathbf{n})}$, then $V\cap H$ is p.g.
@@
$\operatorname{codim}_{V\cap H}V\cap H\cap\operatorname{St}_{\mathbf{s}}=\begin{cases}\operatorname{codim}_{V}V\cap\operatorname{St}_{\mathbf{s}}-1,&\text{if$V\cap\operatorname{St}_{\mathbf{s}}\subset V\cap H$}\\ \operatorname{codim}_{V}V\cap\operatorname{St}_{\mathbf{s}},&\text{otherwise}\end{cases}=\begin{cases}\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})-1,&\text{if $V\cap\operatorname{St}_{\mathbf{s}}\subset V\cap H$}\\ \operatorname{rk}_{P,\mathbf{n}}(\mathbf{s}),&\text{otherwise.}\end{cases}$
To finish, observe that $A\subset\{i:\overline{s}_{i}=n_{i}\}$ if and only if $V\cap\operatorname{St}_{\mathbf{s}}\subset V\cap H$ for a Zariski-general $H$. ∎
@@
$\dim(V\cdot x)=\operatorname{codim}_{V}(V\cap\operatorname{St}_{\mathbf{s}})=\operatorname{rk}_{P}(\mathbf{s}).$ (*)
@@
Otherwise, suppose $N>1$. If $\operatorname{rk}(\mathbf{s})=d$, then the dimension formula implies $Y_{V}\cap O_{\mathbf{s}}$ is nonempty if and only if $\mathbf{s}=\mathbf{n}$. Otherwise, if $\operatorname{rk}\mathbf{s}<d$, then there is $i\in E$ such that $\overline{s}_{i}<n_{i}$. The projection $Y_{V}\to Y_{\pi_{E\setminus i}(V)}$ sends $Y_{V}\cap O_{\mathbf{s}}$ into $Y_{\pi_{E\setminus i}(V)}\cap O_{\pi_{E\setminus i}(\mathbf{s})}$. By 3.9, $\pi_{E\setminus i}(\mathbf{s})$ is not a combinatorial flat of $\pi_{E\setminus i}(V)$, so $Y_{\pi_{E\setminus i}(V)}\cap O_{\pi_{E\setminus i}(\mathbf{s})}$ is empty by induction on $N$. This means $Y_{V}\cap O_{\mathbf{s}}$ is empty too. ∎
@@
We complete the proof for proper combinatorial flats by inducting on $N$. First consider $N=1$, so that $Y_{V}\subset\mathbb{P}^{n_{1}}$. By 4.10, $Y_{V}\cap\mathbb{P}^{n_{1}-1}\subset\mathbb{P}^{\dim V-1}$. Since $Y_{V}\cap\mathbb{P}^{n_{1}-1}$ is a divisor on $Y_{V}$, we must have $Y_{V}\cap\mathbb{P}^{n_{1}-1}=\mathbb{P}^{\dim V-1}$. Since $V$ is polymatroid general, $O_{i}\subset\mathbb{P}^{\dim V-1}$ is a single $V$-orbit, and $\overline{Y_{V}\cap O_{i}}=\mathbb{P}^{i}$ as desired.
@@
Otherwise, suppose $\dim V>1$. For a general hyperplane $H$ in $\mathbb{K}^{\mathbf{n}}$, $Y_{V\cap H}\cap O_{\mathbf{s}}$ is nonempty for every combinatorial flat $\mathbf{s}$ of corank greater than 1 by the induction hypothesis and 3.10. By the same argument, if $F=\{i:s_{i}=n_{i}\}$ is nonempty, then $Y_{V\cap H}\cap O_{\mathbf{s}}$ is nonempty for $H\subset\mathbb{K}^{\mathbf{n}}$ the preimage of a general hyperplane in $\mathbb{K}^{\pi_{F}(\mathbf{n})}$.
@@
where $C$ runs over all irreducible components of $Y_{V}\cap H$. By 4.10, $Y_{V}\cap H\subset\cup_{\mathbf{s}^{\prime}}Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}$, where the union runs over all corank 1 combinatorial flats $\mathbf{s}^{\prime}\leq\mathbf{n}-\mathbf{e}_{i}$. By 4.11, each irreducible component $C$ of $Y_{V}\cap H$ is among those of $\cup_{\mathbf{s}^{\prime}}Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}$, so
@@
with $m_{\mathbf{s}^{\prime}}=0$ if $\dim Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}<d-1$. By 4.11 and 4.6, we may further expand as
@@
Characterization of when $Y_{V}$ intersects $O_{\mathbf{s}}$ is 4.10 and 4.12. 1.3i follows from 4.11, and 1.3iii is obtained by combining 4.11 and 4.6.
@@
Fix $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$. By 3.9, $\mathbf{s}^{\prime}<\mathbf{s}$ is a combinatorial flat of $V$ if and only if $\mathbf{s}_{i}^{\prime}\leq\mathbf{s}_{i}$ for all $i$ and $(\mathbf{s}_{j}^{\prime})_{j:s_{j}=n_{j}}$ is a combinatorial flat of $V^{\prime}:=\pi_{\{j:s_{j}=n_{j}\}}(V)$. Such multisets $\mathbf{s}^{\prime}$ are precisely those that index the nonempty strata of $\overline{Y_{V}\cap O_{\mathbf{s}}}$, since $\overline{Y_{V}\cap O_{\mathbf{s}}}=Y_{V^{\prime}}\times\prod_{s_{i}<n_{i}}\mathbb{P}^{s_{i}}$ by 4.11. 4.11 also implies $O_{\mathbf{s}^{\prime}}\cap(Y_{V^{\prime}}\times\prod_{s_{i}<n_{i}}\mathbb{P}^{s_{i}})=O_{\mathbf{s}^{\prime}}\cap Y_{V}$ whenever both sides are nonempty, which completes the proof. ∎
@@
Theorems 1.3 describes properties of Schubert varieties of subspace arrangements, extending well-known results for Schubert varieties of hyperplane arrangements from [AB16], which correspond to the case $\mathbf{n}=(1,\ldots,1)$. One further such property follows: when $\mathbf{n}=(1,\ldots,1)$, it is well-known that changing $V$ by rescaling each coordinate of $\mathbb{K}^{\mathbf{n}}=\mathbb{K}^{N}$ does not change the isomorphism class of $Y_{V}$. For general $\mathbf{n}$, the isomorphism class of $Y_{V}$ is unchanged if we perform a weighted rescaling of each factor $\mathbb{K}^{n_{i}}$, as described in Section 4.2.
@@
Assume that $V$ realizes a simple polymatroid. The subspaces $V_{ij}=V\cap\operatorname{St}_{j\mathbf{e}_{i}}$ for $1\leq i\leq N$ and $1\leq j\leq n_{i}$ comprise a collection of $N$ flags in $V$. The combinatorial flats record intersection pattern of these flags, as $\operatorname{rk}(\mathbf{s})=\operatorname{codim}_{V}V_{1,s_{1}}\cap V_{2,s_{2}}\cap\cdots\cap V_{N,s_{N}}$. The “polymatroid general” condition on $V$ guarantees that these flags are in general position with respect to one another (see also 3.7).
Example 4.15 (Simplification changes the Schubert variety).
Let $\mathbf{n}=(2,2)$, and consider the subspace $V\subset\mathbb{K}^{\mathbf{n}}=\mathbb{K}^{2}\times\mathbb{K}^{2}$ given by the rowspan of
$\begin{pmatrix}1&0&1&1\\ 0&1&1&-1\end{pmatrix}.$
The data $V\subset\mathbb{K}^{\mathbf{n}}$ realizes the nonsimple polymatroid of 3.18, whose simplification is the boolean matroid on two elements. However, we will see that $Y_{V}$ is not isomorphic to the Schubert variety $(\mathbb{P}^{1})^{2}$ of the simplification. Consider the linear paramitrization $(u,v)\mapsto(u,v,u+v,u-v)$ of $V$, and the corresponding paramatrization of $\iota(V)\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$:
$(u,v)\mapsto[1:u:v+\frac{1}{2}u^{2}]\times[1:u+v:u-v+\frac{1}{2}(u+v)^{2}].$
Recall that we label the coordinates on the right hand side as $[b_{10}:b_{11}:b_{12}]\times[b_{20}:b_{21}:b_{22}]$. The relations between the coordinates $b_{11},b_{12},b_{21},b_{22}$ on the right hand side are generated by the inhomogenious equations
$b_{21}^{2}+4b_{11}-2b_{21}-2b_{22},$
$b_{11}^{2}-2b_{11}-2b_{12}+2b_{21}.$
To take the closure in $\mathbb{P}^{\mathbf{n}}$, we homogenize the above equations in the variables $b_{10},b_{20}$ so that they are bihomogeneous, and then saturate them with respect to the ideal generated by $b_{10}b_{20}$ to obtain the homogenious equations
$160\,b_{11}b_{20}^{2}-32\,b_{12}b_{20}^{2}-96\,b_{10}b_{20}b_{21}+10\,b_{20}b_{21}-b_{21}^{2}+2\,b_{20}b_{22},$
$10\,b_{10}b_{11}b_{20}+b_{11}^{2}b_{20}-2\,b_{10}b_{12}b_{20}-6\,b_{10}^{2}b_{21},$
[…]
$\quad\quad\quad-250\,b_{10}b_{11}b_{21}-25\,b_{11}^{2}b_{21}+50\,b_{10}b_{12}b_{21}+300\,b_{10}^{2}b_{22}.$
One can check (e.g. by using Macaulay2 [GS]) that the subvariety $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ defined by the displayed equations is singular, and therefore not isomorphic to $(\mathbb{P}^{1})^{2}$. $\diamondsuit$
@@
In this section, we establish 1.7, which gives a formula for the cohomology ring of $Y_{V}$. Throughout, we assume that $V$ is polymatroid general. For an independent multiset $\mathbf{b}$, let $c_{\mathbf{b}}>0$ be the coefficient of $[\mathbb{P}^{\mathbf{b}}]$ in $[Y_{V}\cap\mathbb{P}^{\overline{\mathbf{b}}}]$, as in the statement of 1.3.
@@
Let $\kappa:Y_{V}\to\mathbb{P}^{\mathbf{n}}$ denote the inclusion. By 1.3iii, the classes $\{[Y_{V}\cap\mathbb{P}^{\mathbf{s}}]:\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}\}$ are linearly independent in $A_{*}\mathbb{P}^{\mathbf{n}}$, so the pushforward $A_{*}Y_{V}\to A_{*}\mathbb{P}^{\mathbf{n}}$ is injective.
The sets $O_{\mathbf{s}}$ and $Y_{V}\cap O_{\mathbf{s}}$ comprise algebraic cell decompositions of $\mathbb{P}^{\mathbf{n}}$ and $Y_{V}$, respectively. Hence, the Chow groups of these varieties are naturally isomorphic to their Borel-Moore homology groups via the cycle class map [Ful98, Example 19.1.11]. Since $\mathbb{P}^{\mathbf{n}}$ and $Y_{V}$ are compact, their Borel-Moore homology is equal to their singular homology. Summing up, we’ve learned that the singular homology groups of $Y_{V}$ are all free and that the homology pushforward $\kappa_{*}:H_{*}(Y_{V};{\mathbb{Z}})\to H_{*}(\mathbb{P}^{\mathbf{n}};{\mathbb{Z}})$ is injective. Applying the Universal Coefficient Theorem, we learn that $\kappa^{*}:H^{*}(\mathbb{P}^{\mathbf{n}};\mathbb{Q})\to H^{*}(Y_{V};\mathbb{Q})$ is surjective.